OpenStudy (anonymous):

how can i solve a system of equations using substitution x+4y=-15 3x+3y-0

OpenStudy (campbell_st):

for substitution look at the 2nd equation and rewrite it 3x + 3y = 0 or 3x = -3y or x = -y substitute this into the 1st equation (-y) + 4y = -15 or 3y = -15 solve for y when you have y substitute it into either equation to find x

OpenStudy (anonymous):

so how do i solve 3y=-15 do i add 3y to both sides

OpenStudy (campbell_st):

divide both sides of the equation by 3

OpenStudy (anonymous):

ok i got y=5 is that correct

OpenStudy (campbell_st):

not quite it should be y = -5

OpenStudy (campbell_st):

so looking at the 1st equation and substituting y = -5 x + 4*(-5) = -15 so x - 20 = -15 whats x..?

OpenStudy (anonymous):

oh k i got that one so how do i do 9x-8y= -6 x-3y=-7 do i do the same and how

OpenStudy (anonymous):

so i subtract 20 and -15 right.

OpenStudy (anonymous):

i am not sure